# Provably Secure Covert Messaging Using Image-based Diffusion Processes

### v1

@@ IV-E Ciphertext Recovery
1:$s_{\rm key}$: secret key; $\tilde{c_{1}}$: reconstructed ciphertext of first latents; $\tilde{c_{2}}$: reconstructed ciphertext of second latents; ${\rm max\_errs}$: maximum potential errors to correct $\triangleright$ default: 10
@@
6:${\rm err}\leftarrow{\rm Compare}(\tilde{c_{1}},\tilde{c_{2}})$triangleright Identify indexes where the ciphertexts differ.
@@
Once an image has gone through inversion, the receiver has an inverted latents $\tilde{{{\bm{x}}_{T}}}$ (or $(\tilde{{{\bm{x}}_{T}}},\tilde{{{\bm{x}}_{T}}^{\prime}})$ in the case of EDICT). Recall that the receiver also knows ${{\bm{x}}_{T}}$, the initial latents used by the sender for this message. Using ${{\bm{x}}_{T}}$ to identify which component values had magnitudes above the threshold $\tau$, the receiver can compute the ${\rm mask}$ used by the sender and extract the component values associated with it, and then invert the permutation to recover latent components that encode the ciphertext, with redundancy.
@@
This strategy allows us to take into account the magnitude of each component value in the recovery (as opposed to only its sign). For example, it is more likely that inversion error caused $0.02$ to flip to $-0.01$ than it is that $3.0$ flipped to $-3.2$.
@@
Error Correction with EDICT. By contrast, when using EDICT we can perform additional error correction. Our novel proposal works as follows. Due to EDICT we have dual latents $(\tilde{{{\bm{x}}_{T}}},\tilde{{{\bm{x}}_{T}}^{\prime}})$ and therefore we can use Algorithm 2 and Eq. 1 to create two reconstructed ciphertexts $\tilde{c_{1}},\tilde{c_{2}}$, one for each inverted latents. Both ciphertext can be checked to see if the HMAC tag is valid. But if neither is recovered without error, we can use their differences to identify and recover correct bits. If the number of differing bits between $\tilde{c_{1}}$ and $\tilde{c_{2}}$ is small, then we can create a new ciphertext for every combination of differing bits and check if the HMAC tag is valid.
@@ V Formalizing and Analyzing Security
Defining Security & Indistinguishabilities. Recall that the adversary’s proximate goal is to distinguish diffusion images that contain embedded messages from those that do not. Therefore, a straightforward way to define security is as output indistinguishability, i.e., the scheme is secure if the output images produced by the diffusion-based steganographic scheme are indistinguishable from images produced by the diffusion model itself. This is an appealing way to define security, in part because it is easy to evaluate empirically. It suffices to generate a dataset of images produced by the steganographic scheme and a dataset of images produced by the diffusion model alone and compare the two, for example, using an ML-based distinguisher. This is what most related works do [22, 23, 24, 15].
@@
In the rest of this section, we prove the security of our construction. First, we formalize initial latents schemes (Section V-A). We then propose an indistinguishability game for them (Section V-B). In the process, we explain why our proposed approach does not suffer from the same issue as Kim et al. [15] (Section V-C). Finally, we prove our construction’s indistinguishability (Section V-D).
@@ V-C The Problem with Distorting Distributions
Kim et al. [15], our closest prior work, alters the initial latents distribution as part of their embedding process. Indeed, instead of generating a random Gaussian latent space, they use message projection to fill the latent space with values according to the bits they need to embed. So instead of flipping a value’s sign (as we propose), they replace it with 1 for one bits, and -1 for zero bits. Alternatively they pull from a small subset of values, i.e. $\sqrt{2}$ or $-\sqrt{2}$ for 1 bits and 00 for zero bits.
The result is that after embedding their latents do not follow an isotropic Gaussian distribution. Their method can match the mean and standard deviation of the standard Gaussian distribution, but in reality there only three possible values (0, $\sqrt{2}$, or $-\sqrt{2}$). This discrepancy in distribution can be exploited by an adversary to identify covert messages.
Under realistic threat models (such as ours) the adversary does not observe the latents directly. Instead they observe diffusion outputs from which they can attempt to reconstruct the latents by inversion. As we show in experiments (Section VII-A), despite significant loss in the inversion process, the proposal of Kim et al. [15] is not secure. An adversary that approximately reconstructs the latent and trains a simple ML-based distinguisher (e.g., a Random Forest classifier) readily achieves 100% accuracy on distinguishing between their latent spaces and non-steganographic latent spaces.
Interestingly, despite altering the latents distribution, the construction of Kim et al. [15] produces high quality diffusion outputs (as we confirmed experimentally — Section VII-B). The pitfall of their approach is the (implicit) assumption that high output quality implies security.
@@ V-D Achieving Latents Indistinguishability
Let ${{\bm{x}}_{T}}$ be a vector of $k$ random variables each distributed according to a symmetric distribution $\mathcal{D}$, i.e., ${{x}_{T,{i}}}\sim\mathcal{D}$ for $i=1,2,\ldots,k$ i.i.d. Also, let ${\rm Emb}$ denote the transformation of ${{\bm{x}}_{T}}$ into ${{\bm{x}}_{T}}^{\prime}$ in Algorithm 1. Then the distribution of ${{\bm{x}}_{T}}$ is not altered by ${\rm Emb}$, i.e., we have: ${\rm Emb}({{x}_{T,{i}}})\sim\mathcal{D}$ for $i=1,2,\ldots,k$.
@@ VI Experimental Setup
As our LDM, we use CompVis/stable-diffusion-v1-4 [37] with one of two schedulers, DDIM or EDICT. For prompts we use one of ImageNet’s 1000 classes [38]. For the diffusion process, we use 50 inference steps, and we generate 512x512 images. Unless otherwise stated we use a threshold of 0.3 and a redundancy of 6, and the EDICT scheduler. All experiments are done on servers with 12 GB of RAM for the CPU, and an NVIDIA A100 NVSWITCH GPU with 80 GB of memory.
Figure 3: ROC curves and AUCROC for Random Forest classifiers to distinguish latent space with and without embedding. Our method does not alter the latent space and so classifier performance is random guessing. Kim et al. [15] (message projection) do alter the distribution, which leads to (almost) perfect classification.
@@ VII Security Experiments
Although we proved the security of our construction, we can nevertheless evaluate security empirically. For this, we can train ML-based distinguishers to identify any distortions caused by the embedding of our messages. We can also evaluate cover (i.e., output) image quality to determine whether embedding covert messages alters the output distribution in a way that an adversary can detect. For this, we use the Fréchet Inception Distance (FID) [39], CLIP-score [40], and image classifiers as distinguishers.
@@ VII-A Security of the Latent Space
We train Random Forest binary classifiers to distinguish latent spaces. Concretely, we generate 3 sets of 500 latent spaces gathered by inverting a cover image. All images are generated and inverted with a DDIM scheduler. The first set does not contain a steganographic message. The second set contains a message embedded using our method, and the final uses a message embedded by (our re-implementation of) the method used by Kim et al. [15]. Recall that their embedding method overrides the initial latents by assigning 1 bits to either $\sqrt{2}$ or $-\sqrt{2}$ and 0 bits to 0.
@@
Figure 4: QQ plots of the two embedding methods against the natural recovered latent space distribution. Although the graph on the left follows the natural mean and scale well, the clustering around the values inserted into the feature values is easy to detect.
@@ VII-B Security of the cover
We calculate FID using the Pytorch FID implementation [41]. We generate a large set of images. 10,000 without a message embedded, and 10,000 with a message embedded. To establish a base FID, we split the natural dataset in half, then calculated the FID of the two halves. This returned an FID of 13.688. We then calculated the FID of images with a message embedded within them against images without a message, and got an FID of 13.642.
We also gathered the CLIP-scores [40] of the generated images against the prompt used to generate them. Images without a message had an average CLIP-score of $0.363\pm 0.001$. Images with a message had an average score of $0.3628\pm.0008$. Using a t-test with the null hypothesis that these originate from the same distribution, returns a $p$-value of $0.1199$, meaning we do not reject the null hypothesis, and these are likely the same distribution.
@@
Wei et al.[24] 1 1 0.548 $\times$
Peng et al.[32] 0.4 0.995 Unreported $\times$
Kim et al.[15] 1 0.954 Pe=0.423 $\times$
@@ VII-C Steganalysis
We use these same datasets to train several steganalysis models. Clip [42], Xu-net [43], SiaStegNet [44], and YeNet [45]. For CLIP we use OpenAI’s CLIP model with ViT-B/32 transformer architecture. We create two classes “steganographic” and “non-steganographic”, and attempt to classify images according to them. We use the dataset we created in Section VII-B 10,000 images without a message embedded, and 10,000 with a message embedded. Note, this is different from Clip-Score where we compared the image against the prompt used to generate them. Here the model is used to extract features on for a ML-based distinguisher. Our CLIP model uses an Adam optimizer and measures the cross entropy loss. We train for 30 epochs. The model is unable to distinguish between classes.
@@ VIII-B Parameter Tradeoffs
We conducted a grid search over threshold and redundancy levels, shown in Figs. 6(b) and 6(a), to identify the optimal parameter values. We tested thresholds from 00 to $1.5$ incrementing by 0.1, and redundancy from 1 to 16 incrementing by 1. We generated 10 messages at each combination of threshold and redundancy. We then calculated expected bits received by multiplying the number of message bits contained in a single image at those parameter values by the probability of the message successfully being received.
@@ IX Robustness
It is common for image hosting platforms to resize, crop, compress, or re-encode images. Therefore, it is desirable to consider robustness, which is the ability to recover the message despite distortions to the image. A common problem with steganographic methods that embed into the cover image directly, is that they are very susceptible to distortions [46]. When one is attempting to store information inside the least significant bits of a pixel, even minor distortions to the pixel values can undermine reliability. Our premise in this work is that by embedding bits in the initial latents, the deleterious effects of distortions will be much less pronounced.
@@
This quantifies how much the transformation affects the latent space, which directly influences our ability to recover the embedded message. To show the effect of the transformation, we also measure SSIM [47] between the image before and after the transformation.
To further illustrate the advantage of embedding information in the initial latents, we compare our method against two alternatives. The first alternative is to embed the message in the last latent step ${{\bm{x}}_{1}}$ before the latent representation is sent through the encoder to become the full size image. This, theoretically, has the advantage of reducing the inversion error since there is little to no noise added between the embedding of the message and the retrieval. However, this process severely impacts the image quality as can be seen in  Fig. 7. The second alternative is to use a method that embeds bits in the image itself, such as Least Significant Bit (LSB) steganography [46].

### v2

@@ IV-E Ciphertext Recovery
[…]
@@
[…]
@@
[…]
@@
This strategy allows us to take into account the magnitude of each component value in the recovery (as opposed to only its sign). For example, it is more likely that inversion error caused $0.02$ to flip to 0.01-0.01 than it is that $3.0$ flipped to 3.2-3.2.
@@
[…]
@@ V Formalizing and Analyzing Security
Defining Security & Indistinguishabilities. Recall that the adversary’s proximate goal is to distinguish diffusion images that contain embedded messages from those that do not. Therefore, a straightforward way to define security is as output indistinguishability, i.e., the scheme is secure if the output images produced by the diffusion-based steganographic scheme are indistinguishable from images produced by the diffusion model itself. This is an appealing way to define security, in part because it is easy to evaluate empirically. It suffices to generate a dataset of images produced by the steganographic scheme and a dataset of images produced by the diffusion model alone and compare the two, for example, using an ML-based distinguisher. This is what most related works do [peng2023stegaddpm, yu2024cross, wei2023generative, kim2023diffusion].
@@
In the rest of this section, we prove the security of our construction. First, we formalize initial latents schemes (Section V-A). We then propose an indistinguishability game for them (Section V-B). In the process, we explain why our proposed approach does not suffer from the same issue as Kim et al. [kim2023diffusion] (Section V-C). Finally, we prove our construction’s indistinguishability (Section V-D).
@@ V-C The Problem with Distorting Distributions
Kim et al. [kim2023diffusion], our closest prior work, alters the initial latents distribution as part of their embedding process. Indeed, instead of generating a random Gaussian latent space, they use message projection to fill the latent space with values according to the bits they need to embed. So instead of flipping a value’s sign (as we propose), they replace it with 1 for one bits, and -1 for zero bits. Alternatively they pull from a small subset of values, i.e. $\sqrt{2}$ or $\sqrt{2}$ for 1 bits and 0 for zero bits.
The result is that after embedding their latents do not follow an isotropic Gaussian distribution. Their method can match the mean and standard deviation of the standard Gaussian distribution, but in reality there only three possible values (0, $\sqrt{2}$, or $\sqrt{2}$). This discrepancy in distribution can be exploited by an adversary to identify covert messages.
Under realistic threat models (such as ours) the adversary does not observe the latents directly. Instead they observe diffusion outputs from which they can attempt to reconstruct the latents by inversion. As we show in experiments (Section VII-A), despite significant loss in the inversion process, the proposal of Kim et al. [kim2023diffusion] is not secure. An adversary that approximately reconstructs the latent and trains a simple ML-based distinguisher (e.g., a Random Forest classifier) readily achieves 100% accuracy on distinguishing between their latent spaces and non-steganographic latent spaces.
Interestingly, despite altering the latents distribution, the construction of Kim et al. [kim2023diffusion] produces high quality diffusion outputs (as we confirmed experimentally — Section VII-B). The pitfall of their approach is the (implicit) assumption that high output quality implies security.
@@ V-D Achieving Latents Indistinguishability
Let ${\bm{x}}_{T}$} be a vector of $k$ random variables each distributed according to a symmetric distribution $\mathcal{D}$, i.e., ${{x}_{T,{i}}}\sim\mathcal{D}$ for $1,2,\ldots,k i.i$.d. Also, let $\rm Emb$} denote the transformation of ${\bm{x}}_{T}$} into ${\bm{x}}_{T}$$^{\prime}$ in Algorithm 1. Then the distribution of ${\bm{x}}_{T}$} is not altered by $\rm Emb$}, i.e., we have: ${\rm Emb}({{x}_{T,{i}}})\sim\mathcal{D}$ for $i=1,2,\ldots,k$.
@@ VI Experimental Setup
As our LDM, we use CompVis/stable-diffusion-v1-4 [Rombach_2022_CVPR] with one of two schedulers, DDIM or EDICT. For prompts we use one of ImageNet’s 1000 classes [deng2009imagenet]. For the diffusion process, we use 50 inference steps, and we generate 512x512 images. Unless otherwise stated we use a threshold of 0.3 and a redundancy of 6, and the EDICT scheduler. All experiments are done on servers with 12 GB of RAM for the CPU, and an NVIDIA A100 NVSWITCH GPU with 80 GB of memory.
Figure 3: ROC curves and AUCROC for Random Forest classifiers to distinguish latent space with and without embedding. Our method does not alter the latent space and so classifier performance is random guessing. Kim et al. [kim2023diffusion] (message projection) do alter the distribution, which leads to (almost) perfect classification.
@@ VII Security Experiments
Although we proved the security of our construction, we can nevertheless evaluate security empirically. For this, we can train ML-based distinguishers to identify any distortions caused by the embedding of our messages. We can also evaluate cover (i.e., output) image quality to determine whether embedding covert messages alters the output distribution in a way that an adversary can detect. For this, we use the Fréchet Inception Distance (FID) [heusel2017gans], CLIP-score [hessel2021clipscore], and image classifiers as distinguishers.
@@ VII-A Security of the Latent Space
We train Random Forest binary classifiers to distinguish latent spaces. Concretely, we generate 3 sets of 500 latent spaces gathered by inverting a cover image. All images are generated and inverted with a DDIM scheduler. The first set does not contain a steganographic message. The second set contains a message embedded using our method, and the final uses a message embedded by (our re-implementation of) the method used by Kim et al. [kim2023diffusion]. Recall that their embedding method overrides the initial latents by assigning 1 bits to either $\sqrt{2}$ or $\sqrt{2}$ and 0 bits to 0.
@@
Figure 4: $Q$ plots of the two embedding methods against the natural recovered latent space distribution. Although the graph on the left follows the natural mean and scale well, the clustering around the values inserted into the feature values is easy to detect.
@@ VII-B Security of the cover
We calculate FID using the Pytorch FID implementation [Seitzer2020FID]. We generate a large set of images. 10,000 without a message embedded, and 10,000 with a message embedded. To establish a base FID, we split the natural dataset in half, then calculated the FID of the two halves. This returned an FID of 13.688. We then calculated the FID of images with a message embedded within them against images without a message, and got an FID of 13.642.
We also gathered the CLIP-scores [hessel2021clipscore] of the generated images against the prompt used to generate them. Images without a message had an average CLIP-score of $0.363\pm 0.001$. Images with a message had an average score of $0.3628\pm.0008$. Using a t-test with the null hypothesis that these originate from the same distribution, returns a $p$-value of $0.1199$, meaning we do not reject the null hypothesis, and these are likely the same distribution.
@@
Wei et al.[wei2023generative] 1 1 0.548 $\times$
Peng et al.[peng2024ldstega] 0.4 0.995 Unreported $\times$
Kim et al.[kim2023diffusion] 1 0.954 Pe=0.423 $\times$
@@ VII-C Steganalysis
We use these same datasets to train several steganalysis models. Clip [radford2021learning], Xu-net [xu2016structural], SiaStegNet [you2020siamese], and YeNet [ye2017deep]. For CLIP we use OpenAI’s CLIP model with ViT-B/32 transformer architecture. We create two classes “steganographic” and “non-steganographic”, and attempt to classify images according to them. We use the dataset we created in Section VII-B 10,000 images without a message embedded, and 10,000 with a message embedded. Note, this is different from Clip-Score where we compared the image against the prompt used to generate them. Here the model is used to extract features on for a ML-based distinguisher. Our CLIP model uses an Adam optimizer and measures the cross entropy loss. We train for 30 epochs. The model is unable to distinguish between classes.
@@ VIII-B Parameter Tradeoffs
We conducted a grid search over threshold and redundancy levels, shown in Figs. 6(b) and 6(a), to identify the optimal parameter values. We tested thresholds from 0 to $1.5$ incrementing by 0.1, and redundancy from 1 to 16 incrementing by 1. We generated 10 messages at each combination of threshold and redundancy. We then calculated expected bits received by multiplying the number of message bits contained in a single image at those parameter values by the probability of the message successfully being received.
@@ IX Robustness
It is common for image hosting platforms to resize, crop, compress, or re-encode images. Therefore, it is desirable to consider robustness, which is the ability to recover the message despite distortions to the image. A common problem with steganographic methods that embed into the cover image directly, is that they are very susceptible to distortions [petitcolas1999information]. When one is attempting to store information inside the least significant bits of a pixel, even minor distortions to the pixel values can undermine reliability. Our premise in this work is that by embedding bits in the initial latents, the deleterious effects of distortions will be much less pronounced.
@@
This quantifies how much the transformation affects the latent space, which directly influences our ability to recover the embedded message. To show the effect of the transformation, we also measure SSIM [wang2004image] between the image before and after the transformation.
To further illustrate the advantage of embedding information in the initial latents, we compare our method against two alternatives. The first alternative is to embed the message in the last latent step ${\bm{x}}_{1}$} before the latent representation is sent through the encoder to become the full size image. This, theoretically, has the advantage of reducing the inversion error since there is little to no noise added between the embedding of the message and the retrieval. However, this process severely impacts the image quality as can be seen in  Fig. 7. The second alternative is to use a method that embeds bits in the image itself, such as Least Significant Bit (LSB) steganography [petitcolas1999information].